# Large-scale Long-tailed Disease Diagnosis on Radiology Images

### v2

@@ 4.1 Problem Formulation
the model is composed of a visual encoder $\Phi_{\text{visual}}$, a fusion module $\Phi_{\text{fuse}}$ and a classifier $\Phi_{\text{cls}}$.
4.2 Architecture
Figure 6: The architecture of our proposed visual encoder and fusion module, together with the knowledge enhancement strategy. (a) shows the details of the vision encoder. We design two variants to fit in the two main visual backbones, i.e., ResNet and ViT. (b) shows the transformer-based fusion module, enabling case-level information fusion. (c) shows the knowledge enhancement strategy. We first pre-train a text encoder with extra medical knowledge with contrastive learning, i.e., synonyms, descriptions and hierarchy, termed as knowledge encoder and then we view the text embedding as a natural classifier to guide the diagnosis classification.
Our proposed architecture consists of two key components: (i) a visual encoder that processes 2D or 3D input scan; (ii) a transformer-based fusion module, merging all information to perform case-level diagnosis.
4.2.1 Visual Encoder
We consider two popular variants of the visual encoder, namely, ResNet [30] and ViT [23], as shown in Figure 6 (a). The visual encoding progress can generally be formulated as:
$v_{i}=\Phi_{\text{visual}}(x_{i})\in\mathcal{R}^{d}.$ (2)
where $x_{i}\in\mathcal{R}^{C\times H\times W\times(D)}$ denotes the input scan, $C,H,W$ refer to the image channel, height and width of the input scan respectively. $D$ is optional, and only available for 3D input scans. The main challenge for architecture design comes from the requirement to process scans in both 2D and 3D formats. Here, we train separate normalisation modules to convert the 2D or 3D inputs into feature maps of same resolution, and further passed into shared encoding module.
ResNet. For 3D scans, they are first fed into 3D ResNet, followed by average pooling on depth to aggregate the information of the extra dimension; while for 2D scans, they are fed to the 2D ResNet to perform the same down-sampling ratio on height and width as for 3D. After normalization, both 2D and 3D scans are transformed into feature maps with same resolution, $f_{i}\in\mathcal{R}^{d_{\text{res}}\times h\times w}$, where $d_{\text{res}}$ is the intermediate feature dimension and $h,w$ denote the normalized size. Then, the feature maps are passed into a shared ResNet to get the final visual embedding.
ViT. For 3D scans, we convert the input volume into a series of non-overlapped 3D cubes, and pass them into MLP projection layers, to get vector embeddings; while for 2D scans, the input scan is broken into 2D patches, and projected into vector embeddings with another set of MLP layers. As ViT enables to handle sequences of variable tokens, we can now pass the resulting vectors into a shared ViT to get the final visual embedding. For position embedding, we adopt two sets of learnable position embeddings, one for 2D input, the other for 3D, further indicating the network about the input dimenstion.
4.2.2 Fusion Module
For case-level diagnosis, we propose to aggregate information from multiple scans with a trainable module. As shown in Figure 6 (b), we adopt transformer encoders, specifically, we first initialize a set of learnable modality embeddings, denoted as $\{p_{1},\dots,p_{M}\}$, where $M$ denotes the total number of possible imaging modalities. Given certain visual embedding ($v_{i}$) from modality $j$, we first add the corresponding modality embedding ($p_{j}$) to it, indicating which radiologic modality it is from. Then we feed all visual embeddings from one case into the fusion module, and output the fused visual embedding $v_{\text{fuse}}\in\mathcal{R}^{d}$, from the “[cls]” token, similar to paper [23], denoted as:
$v_{\text{fuse}}=\Phi_{\text{fuse}}(\Phi_{\text{visual}}(x_{1}),\Phi_{\text{visual}}(x_{2}),\cdots,\Phi_{\text{visual}}(x_{S}))\in\mathcal{R}^{d}.$ (3)
Till here, we have computed the case-level visual embedding, which can be passed into classifier for disease diagnosis. Here, we adopt a knowledge-enhanced training strategy [79, 84], that has shown to be superior in long-tailed recognition problem, as detailed in the following section.
4.3 Knowledge-enhanced Training
With knowledge enhancement, we hope to leverage the rich domain knowledge in medicine to enhance the long-tailed classification. Our key insight is that the long-tailed diseases may fundamentally have some shared symptoms or radiologic pathologies with the head classes, which could be encoded in text format.
In detail, we first pre-train a text encoder with medical knowledge, termed as knowledge encoder, where names of similar disorder or diseases are projected to similar embeddings, for example, ‘lung disease’ is closer to ‘pneumonia’ in the embedding space than ‘brain disease’. Then we freeze the text embeddings, and use them to guide the training of vision encoder, as shown in Figure 6 (c).
Knowledge Encoder Pre-training. We leverage several extra knowledge bases to pre-train a knowledge encoder, including Radiopaedia, ICD10-CM, and UMLS. Specifically, for each disorder term, we collect its definitions, radiologic features from Radiopaedia articles, synonyms, clinical information, and hierarchy structure from ICD10-CM, as well as definitions from UMLS. We aim to train the text encoder with the following considerations:
  • Synonyms. If two terms are identified as similar synonyms, we expect them to be close in text embedding space, like ‘Salmonella enteritis’ and ‘Salmonella gastroenteritis’.
  • Hierarchy. In the context of hierarchical relationships, we expect that if a disease is a fine-grained classification of another disease, its embeddings should be closer than those of unrelated diseases. For example, ‘J93, Pneumothorax and air leak’ should exhibit closer embeddings with ‘J93.0, Spontaneous tension pneumothorax’ than with ‘J96.0, Acute respiratory failure’.
  • Descriptions. For terms associated with descriptions, radiologic features or clinical information, we expect their embeddings to be close in the text embedding space. For example, ‘Intestinal Neuroendocrine Tumor’ and ‘A well-differentiated, low or intermediate grade tumor with neuroendocrine differentiation that arises from the small or large intestine’.
We start from an off-the-shelf text encoder, namely, MedCPT-Query-Encoder [37], and adopt contrastive learning for further finetuning. Given a target medical terminology name encoded by the text encoder, denoted as $f_{\text{tar}}$, its corresponding medical texts, i.e., synonyms, containing terminologies or related descriptions are treated as positive cases. Similarly, we encode them with the text encoder, denoted as $f^{+}$, and other non-related text embedding is treated as negative cases $f^{-}$. Note that, we keep positive and negative cases consistent in format, e.g., when the positive case is related to description sentence, the negative cases are always some non-related description sentences instead of the short name words. The final objective can be formulated as:
where $\tau$ refers to the temperature, and $N$ denotes total sampled negative cases. By optimizing the contrastive loss we can further finetune the text encoder, resulting in a knowledge encoder, termed as $\Phi_{k}$
$t_{j}=\Phi_{k}(T_{j})\in\mathcal{R}^{d},$ (5)
where $t_{j}$ denotes the text embedding. The resulting text embeddings are used for case-level diagnosis:
$p=v_{\text{fuse}}\cdot t\in\mathcal{R}^{c},$ (6)
5 Experiments
In this section, we will introduce our evaluation settings. Specifically, we first establish a benchmark for case-level multi-modal, multi-scan, long-tailed disorder/disease diagnosis. Second, we treat RP3D-DiagDS as a large-scale dataset for pre-training, and evaluate its transferring ability to various existing datasets.
5.1 Long-tailed Classification on RP3D-DiagDS
With the proposed dataset, we consider the problem as a multi-label classification task under a long-tailed distribution. In this section, we provide an overview of the training and evaluation protocols employed in our study. We first introduce the dataset split, followed by details of evaluation metrics for assessment.
5.1.1 Dataset Split
We randomly split our dataset into training (train and validation) and test sets in a (7:1):2 ratio. We treat the class with [100:][100:][ 100 : ] positive cases as head classes, $[30,100)$ cases as medium classes, and [:30)[:30)[ : 30 ) cases as tail classes. Consequently, from a total of 38,858 cases, the training set comprises 27,201 cases, and the test set includes 7,772 cases. Following on the dataset construction procedure, we perform classification task on two class sets at different granularities: (i) 5568 disorders + normal; (ii) 930 ICD-10-CM codes + normal. On different sets, we will have different head/medium/tail class set following our definition. Note that, the normal class is always treated as one of the head classes. Next, we will only talk about the abnormal classes.
Disorders. At the disorder level, there are 85 head classes, 470 medium classes, and 5014 tail classes. Notably, as the case number of some tailed classes is extremely small, some classes may only appear in training or test sets, but not both. This issue only happens to the tail classes. It is important to note that our knowledge encoder enables to get embedding for unseen diseases, ensuring that evaluation is not affected by such challenge. As a result, there are 5,305/3,399 classes in our training and testing split, respectively.
ICD-10-CM. At ICD-10-CM level, disorders with similar meanings have been mapped to same codes, for example, ‘tuberculosis’, ‘primary pulmonary tuberculosis’ all correspond to ‘A15 respiratory tuberculosis’. This merge operation results in more cases in one class, consequently more head classes. Specifically, we have 165 head classes, 230 medium classes, and 536 tail classes. Similarly, some tail classes may only be in training or testing split, resulting in 902/759 classes for our final training and testing splits, respectively.
It is important to point that while we ensure the inclusion of head and medium classes in both the training and test sets, some tail classes are inevitably exclusive from certain divisions, which is treated as our future work to collect more samples of such disorders. We report the results separately for head/medium/tail classes, while focusing primarily on the head classes.
5.1.2 Evaluation Metrics
In this section, we describe the evaluation metrics in detail. Note that, the following metrics can all be calculated per class. For multi-class cases, we all use macro-average on classes to report the scores by default. For example, “AUC” for multi-class classification denotes the “Macro-averaged AUC on classes”.
F1 and MCC. F1 score is the harmonic mean of the precision and recall. It is widely used in diagnosis tasks. MCC [18] denotes the Matthews Correlation Coefficient metric. It ranges from $-1$ to $1$ and can be calculated as follow:
Both metrics need a specific threshold to compute, and we choose one by maximizing F1 on the validation set following former papers [59, 79, 80, 84].
5.2 Transfer Learning to External Datasets
In addition to evaluating on our own benchmark, we also consider transferring our final model to other external datasets, demonstrating its transferring abilities on image distribution shift and label space change. In the following sections, we start by introducing the used external datasets, that cover various medical imaging modalities and anatomies. Then we detail the fine-tuning settings.
5.2.1 External Datasets
We choose external evaluation dataset with the following principles:
  • Imaging Modalities. We hope to cover most radiologic modalities in our external evaluation, e.g., 2D X-ray, 2D CT/MRI slices and 3D CT/MRI scans, demonstrating our model can benefit all of them;
  • Human Anatomies. We hope to cover many human anatomies, e.g., brain, head and neck, chest, spine abdomen and limb, demonstrating our model can benefit all of them;
As a result, we pick the following datasets, and report AUC scores to compare with others on these external evaluation.
  • ADNI [35] is a 3D brain MRI dataset focused on Alzheimer’s disease comprising 112141 images, including AD, MCI, CN and other classifications. We random split it follow 63,846/15,962 for training and testing and reproduce the state-of-the-art method on it.
  • MosMedData [57] is a 3D chest CT dataset on 5-level COVID-19 grading comprising 1,110 images. We follow the official split. We random split it follow 888/222 for training and testing and reproduce the state-of-the-art method on it.
5.2.2 Fine-tuning Diagnosis
Our final model can serve as a pre-trained model and be fine-tuned on each downstream task to improve the final performance, demonstrating the merit of our dataset. Specifically, for the dataset with single image input, we simply adopt the pre-trained visual encoder module, i.e., discarding the fusion module. While for the datset with multi image input, we will use them all. In both cases, the final classification layer will be trained from scratch. In addition to using all available external training data, we also consider to use 1%, 10%, 30% portion data for few-shot learning.
5.3 Implementation Details
At training time, we consider two diagnose tasks in different granularities: Disorder-level classification (5569 classes) and ICD10-level classification (931 Classes). The image input will all be resized to $512\times 512\times D$ in height, width and depth respectively. For 3D scans, the depth is treated as a factor for ablation study, and will be discussed in experiment section. In vision encoder, we adopt two separated modules for normalization and a shared module to compute the final embedding. The detailed architecture design will also be further discussed in our ablation study. In fusion module, we use a 6-layer transformer encoder with learnable ‘[cls]’ token for final prediction.
For augmentation, we adopt Gaussian Noise, Contrast Adjustment, Affine Variation, and Elastic Deformation, implemented from the MONAI [16] package. For optimization, we utilize the AdamW optimizer with a cosine learning rate curve, setting the maximum learning rate to $lr=1\times e^{-5}$, with an adjustable batch size ranging from 4 to 32 depending on the input image depth and model scale to avoid out-of-memory error. The total training duration spans 100 epochs, with the initial 5 epochs for warm-up. At fine-tuning stage, we adopt the similar model architecture and optimization setting.
6 Results
We conduct an ablation study to identify the optimal architecture and hyper-parameters for our model. Then, we present the evaluation results, focusing on the ICD-10-CM classification and disorders classification. Lastly, we use our dataset for large-scale pre-training, and finetune it on various external datasets, to demonstrate the model’s transferring ability.
6.1 Ablation Study
To explore the optimal model architecture and parameter configurations, including the fusion strategy, visual encoder architecture, the depth of 3D scans, and augmentation, we conduct a series of ablation studies on a subset of original dataset, comprising 200 disorder categories with most cases, termed as SubSet@200. In the default experiment setting, we use 16 as the 3D scan depth, and ResNet as visual backbone, without knowledge enhancement and augmentation strategy. While conducting ablation study on certain component, we keep other setting unchanged.
6.1.1 Visual Encoder Architecture
In this section, we investigate different backbone architectures for diagnosis, specifically, we compare the ResNet-based and the ViT-based models. As shown in Table 2, “Seperated” denotes the separated visual encoder to perform 2D and 3D normalization and “Shared” denotes the shared encoder part for both 2D and 3D scans. We make two observations, (i) ResNet-based architecture performs better than ViT-based ones, (ii) increasing the capacity of shared encoder is more beneficial, e.g., ResNet-34+ResNet-18 vs. ResNet-18+ResNet-34. (iii) Deeper ResNet structure has very limited improvement, e.g., ResNet-34+ResNet-34 vs. ResNet-18+ResNet-34.
@@ 6.1.1 Visual Encoder Architecture
6.1.2 Image Dimension
Here, we aim to investigate the effect of input resolution, i.e., increasing the depth of input volumes. Due to the constraints of GPU physical memory, we experiment with 16, 24, and 32 as the depths for 3D input volume. We employ trilinear interpolation to resample 3D scans to the same size.
Table 3 illustrates the performance by varying the depths of 3D images. It can be observed that an increase in depth can bring clear performance gain, suggesting that the detailed depth information is critical to perform diagnosis. Consequently, with more slices, the model tend to yield more favorable results.
Table 3: Ablation study on different 3D scan depth size settings. Considering the computational cost, we evaluated the commonly seen 16, 24, 32 depth sizes on our datsets.
3D Image Depth AUC AP F1 MCC R@0.01 R@0.05 R@0.1
16 87.06 11.27 17.36 19.23 21.48 44.38 61.54
24 87.49 11.41 17.84 19.39 21.76 45.72 62.53
32 88.13 11.98 18.02 19.67 22.84 47.46 63.91
6.1.3 Augmentation
Here, we evaluate the effectiveness of augmentation on our dataset. Specifically, we adopt four augmentation strategies with $15\%$ probability each, namely, Gaussian Noise, Contrast Adjustment, Affine Variation, and Elastic Deformation. As a result, about half of the training data will be applied at least one augmentation in each training batch. As shown in Table 4, adopting data augmentation shows notable performance improvement. This improvement is particularly evident on AUC and AP when employing the ViT as the visual backbone while, still, ResNet-based model performs better.
Table 4: Ablation on adopting augmentation strategy. In the table, we show the effectiveness of adopting reasonable data augmentation for our task. “ResNet” denotes the ResNet-based visual encoder architecture and “ViT” denotes the ViT-basde one.
Augmentation Visual Encoder AUC AP F1 MCC R@F0.01 R@F0.05 R@F0.1
YES ResNet 88.96 13.92 22.04 23.08 25.75 51.61 66.47
ViT 84.22 8.69 16.36 17.18 19.83 38.70 54.04
NO ResNet 87.06 11.27 17.36 19.23 21.48 44.38 61.54
ViT 81.69 6.40 14.71 15.30 18.11 34.73 48.84
6.1.4 Summary
In conclusion, we find the ResNet-based model with augmentation strategy and unifying the 3D scan depth to 32 are the most suitable setting for our long-tailed case-level multi-modal diagnosis task, which will be used for training on the entire dataset in the following sections.
6.2 Evaluation on Rad3D-DiagDS
In this section, we train our model on the entire Rad3D-DiagDS training set, and evaluate on the test split respectively, as shown in Table 5, and the AUC curve comparison is shown in Figure 7.
Specifically, we carry out the experiments at two levels, i.e., disorder and ICD-10-CM classes. We denote fusion module as FM for short and knowledge enhancement as KE. In cases where without FM or KE, we adopt max pooling on the predictions of different images from the same case (Check supplementary for more details), serving as a baseline. Then, we add the fusion module and knowledge enhancement step-by-step, to improve the model’s performance.
@@ 6.2 Evaluation on Rad3D-DiagDS
Table 5: Classification results on Disorders and ICD-10-CM levels. In the table “FM” represents the fusion module and “KE” represents the knowledge enhancement strategy. We report the results on Head/Medium/Tail class sets separately.
Granularity Class Methods Metrics
FM
KE
AUC AP F1 MCC R@0.01 R@0.05 R@0.1
Disorders Head
91.91 13.56 22.08 23.66 30.88 60.50 74.70
93.56 14.2 24.77 25.84 36.19 65.21 79.83
94.24 15.33 25.68 26.71 37.06 66.55 81.37
Medium
93.62 9.19 17.16 20.42 28.24 59.32 73.80
94.03 11.47 19.18 23.14 30.04 63.89 78.69
94.69 12.38 20.64 24.07 31.52 65.34 78.73
Tail
88.55 4.34 8.03 12.96 9.13 27.46 41.43
90.01 8.97 9.35 13.30 10.04 27.02 42.32
90.64 9.25 9.89 14.38 10.98 27.98 43.53
ICD-10-CM Head
88.29 12.01 19.80 20.85 22.95 50.12 62.25
89.55 13.89 21.92 23.41 25.04 52.19 67.26
90.89 14.37 22.67 24.29 26.11 53.82 69.16
Medium
88.37 7.85 15.58 18.02 22.88 49.62 65.25
89.90 8.98 17.33 19.89 24.37 51.33 64.93
91.67 9.56 18.32 20.77 25.68 52.85 66.63
Tail
85.28 4.02 7.53 11.71 7.94 21.28 34.41
86.34 4.77 8.81 12.85 8.83 22.36 37.11
86.55 4.81 8.69 12.74 8.86 22.75 37.97
(a) ROCs on Disorders (Head).
(b) ROCs on Disorders (Medium).
(c) ROCs on Disorders (Tail).
(d) ROCs on ICD-10-CM (Head).
(e) ROCs on ICD-10-CM (Medium).
(f) ROCs on ICD-10-CM (Tail).
Figure 7: ROC curves on Disorders and ICD-10-CM, including head/medium/tail parts respectively. The shadow in the figure shown the 95% CI (Confidence interval) and FM, KE are short for Fusion Module and Knowledge Enhancement.
@@
Discussion on Fusion Module. To demonstrate the effectiveness of fusion module, we start with a baseline without FM or KE. As shown in Table 5, adding the fusion module can greatly improve the results on Head, Medium and Tail classes at both disorder and ICD-10-CM levels, showing the critical role for case-level information fusion in diagnosis task. These results align well with our expectations. In clinical practice, the examination of one modality for a diagnosis is often insufficient. A thorough and meticulous diagnostic process typically involves an integrated review of all test results. Each test is weighted differently, depending on how its results correspond with other tests. Our fusion model adeptly mirrors this comprehensive approach, demonstrating its effectiveness in simulating the nuanced process of clinical diagnosis.
@@
Discussion on Knowledge Enhancement. As shown in Table 5, incorporating knowledge enhancement further promotes the final diagnosis performance on both disorder and ICD-10-CM cases, validating the assumption that a better text embedding space trained by rich domain knowledge can greatly help the visual representation learning.

### v3

@@ 2 RESULTS
where the model is composed of a visual encoder $\Phi_{\text{visual}}$, a fusion module $\Phi_{\text{fuse}}$ and a classifier $\Phi_{\text{cls}}$.
RP3D-DiagDS Statistics
Here, we provide a detailed statistical analysis of our collected dataset, namely RP3D-DiagDS, from three aspects: modality coverage, anatomy coverage, and disease coverage. As shown in Supplementary Section E, compared with existing datasets, our dataset has great advantages in terms of anatomy richness, modal richness, and the number of images and categories. Furthermore, we split the dataset based on the number of cases on different levels.
Analysis on Modality Coverage. RP3D-DiagDS comprises images from 9 modalities, namely, computed tomography (CT), magnetic resonance imaging (MRI), X-ray, Ultrasound, Fluoroscopy, Nuclear medicine, Mammography, DSA (angiography), and Barium Enema. Each case may include images from multiple modalities, to ensure the precise and comprehensive diagnosis of disorders. Overall, approximately 19.4% of the cases comprise images from two modalities, while around 2.9% involve images from three to five modalities. The remaining cases are associated with image scans from a single modality. The distribution of modalities among all abnormal and normal samples are illustrated in Figure 2a.
Analysis on Anatomy Coverage. RP3D-DiagDS comprises images from various anatomical regions, including the head and neck, spine, chest, breast, abdomen and pelvis, upper limb, and lower limb, providing comprehensive coverage of the entire human body. The statistics are shown in Figure 2b.
Figure 2: The data distribution analysis on RP3D-DiagDS. a The distribution of imaging modalities of abnormal (left) and normal (right) cases in RP3D-DiagDS. Each label is annotated with the class name, number of cases, and the corresponding proportion. b The distribution of imaging anatomies of abnormal (left) and normal (right) cases in RP3D-DiagDS. c Case distribution on image numbers. In the bar plot, We show the distribution for the number of images in one case. In RP3D-DiagDS, each case may include multiple images from patient history scans, different modalities, and different angles or conditions. d Case distribution on classes. We demonstrate the long-tailed distributions for disorder and ICD-10-CM classes. We also categorize these classes into three categories: “head class”, “body class” and “tail class” based on the number of cases. Notably, to better show the main part of the case distributions, we clip the axes, indicated by the dotted axes lines.
Analysis on Disease Coverage. For both disorder and disease classification, each case can correspond to multiple disorders, resulting in RP3D-DiagDS a multi-label classification dataset. As shown in Figure 2d, the distributions exhibit extremely long-tailed patterns, rendering such 2D & 3D image classification problems as a long-tailed multi-label classification task. Totally there are 5568 disorders mapped into 930 ICD-10-CM classes. We define the “head class” category with case counts greater than 100, the “body class” category with case counts between 30 and 100, and the “tail class” category with case counts less than 30.
Ablation Studies
We first carry out a series of ablation studies to explore the optimal model architecture and parameter configurations, including the fusion strategy, visual encoder architecture, the depth of 3D scans, and augmentation. Note that, considering the original RP3D-DiagDS is too large to carry out extensive empirical experiments, we construct a subset by only keeping the top 200 classes with the most cases, accordingly, tracking the performance on these 200 classes for ablation studies.
As simple experiment baselines, we use images with spatial resolution of 512x512, 16 as the 3D scan depth, and ResNet-18+ResNet-34 as the visual backbone, without knowledge enhancement or augmentation strategy, adopting the basic max pooling fusion strategy introduced in Supplementary Section B. While conducting an ablation study on certain component, we keep other settings unchanged. More results can be found in Supplementary Section C. Herein, we mainly describe the critical results on comparing image encoder architectures.
@@
ResNet-ViT ResNet-34 6-layer ViT 88.74 11.52 17.86 20.05 21.92 44.63 63.09
ResNet-50 6-layer ViT 89.53 11.75 19.59 20.61 23.18 51.34 67.39
ResNet-34 12-layer ViT 88.93 11.49 18.07 20.09 22.38 45.23 65.04
ResNet-50 12-layer ViT 89.56 11.73 19.73 20.11 22.58 51.64 67.92
Specifically, we investigate three backbone architectures variants for realizing unified diagnosis with arbitrary 2D and 3D scans, i.e., ResNet-based, ViT-based and ResNet-ViT-mixing. More details on architecture design can be found in Section 4.2.1. As shown in Table 1, we make three observations, (i) The performance of ResNet-ViT-mixing based model significantly surpasses that of ViT-based and ResNet-based ones. (ii) Enhancing the feature dimension of the vision encoder results in notable improvements, e.g.,ResNet-34 + 6-layerViT vs. ResNet-50 + 6-layerViT. (iii) Increasing the number of shared encoder layers in the ViT-based model does not yield a significant enhancement in performance but incurs substantial computational overhead, e.g., ResNet-50 + 6-layer ViT vs. ResNet-50 + 12-layer ViT.
Based on the comprehensive ablation results shown in Supplementary Section C, we choose the ResNet-ViT-mixing model with augmentation strategy and unifying the 3D scan depth to 32 as the most suitable settings for our following long-tailed case-level multi-modal diagnosis ability evaluation. After deciding the basic configurations, we evaluate the diagnosis ability on both our proposed internal RP3D-DiagDS dataset and various external benchmarks. In internal evaluation, we mainly show the effectiveness of our newly proposed modules, considering the problem as a multi-label classification task under a long-tailed distribution. In external evaluation, we systematically investigate the boundary of our final model, serving as a diagnosis foundation model under both fine-tuning and zero-shot settings.
Evaluation on Internal Benchmark
In this part, our model is trained on the entire Rad3D-DiagDS training set (29536 cases), evaluated on its test split (7772 cases), as detailed in Table 2. The comparison of AUC curves is depicted in Figure 3. We conduct experiments at two levels—disorder and ICD-10-CM classes—to evaluate the model’s performance, both with and without text knowledge enhancement. These evaluations consider two integration strategies: early fusion and late fusion. For comprehensive definitions and applications of these fusion strategies, please see Supplementary Section B for more details. The model’s effectiveness is assessed using various metrics on the test set, including the AUC curve’s confidence interval (CI), to gauge its stability and consistency.
@@
Table 2: Classification results on Disorders and ICD-10-CM levels. In the table “FM” represents the fusion module and “KE” represents the knowledge enhancement strategy. We report the results on Head/Medium/Tail category sets separately.
Granularity Category Methods Metrics
FM KE AUC AP F1 MCC R@0.01 F@0.05 R@0.1
Disorders Head 92.44 15.34 24.95 26.14 34.11 63.70 76.70
93.61 18.16 28.15 29.30 39.89 67.74 80.10
94.41 20.27 30.21 32.27 41.93 71.09 81.15
Medium 93.74 11.12 19.88 22.67 32.62 62.29 73.63
94.79 13.51 22.96 25.80 37.63 66.25 78.06
95.14 15.95 25.82 28.84 42.49 68.73 79.39
Tail 88.10 6.04 10.35 15.06 9.52 22.48 32.94
90.13 6.65 10.98 15.84 11.43 27.37 43.01
90.96 7.75 12.68 17.49 13.13 28.88 44.08
ICD-10-CM Head 89.83 12.31 20.27 21.58 23.81 52.05 68.45
90.25 13.94 21.78 22.98 25.94 55.22 69.15
91.27 14.59 22.81 25.12 27.87 57.75 72.32
Medium 90.58 8.42 16.42 19.41 23.80 50.80 66.42
91.12 9.18 18.02 20.05 27.10 53.77 67.45
92.01 10.34 19.08 22.16 28.81 55.55 69.86
Tail 86.91 4.82 8.64 12.70 10.13 29.82 51.04
87.32 5.09 9.29 13.23 11.21 30.00 49.39
88.11 5.57 10.48 14.68 12.75 30.11 51.77
    *All indicators are presented as percentages, with the % omitted in the table.
Analysis on Fusion Module (FM). To demonstrate the effectiveness of the fusion module, we start with a baseline that adopts max pooling on the predictions of different images from the same case (check Supplementary Section B for more details). Then, we add the fusion module and knowledge enhancement step-by-step, to improve the model’s performance.
@@
As shown in Table 2, adding the fusion module can greatly improve the results on Head, Medium, and Tail classes at both disorder and ICD-10-CM levels, showing the critical role of case-level information fusion in diagnosis tasks. These results align well with our expectations, as in clinical practice, the examination of one modality for a diagnosis is often insufficient. A thorough and meticulous diagnostic process typically involves an integrated review of all test results. Each test is weighted differently, depending on how its results correspond with other tests. Our fusion model adeptly mirrors this comprehensive approach, demonstrating its effectiveness in simulating the nuanced process of clinical diagnosis.
@@
Analysis on Knowledge Enhancement (KE). We experiment with the knowledge-enhanced training strategy, where we first leverage the rich domain knowledge to pre-train a knowledge encoder with natural language and use it to guide the visual representation learning for disease diagnosis. As shown in Table 2, knowledge enhancement further promotes the final diagnosis performance on both disorder and ICD-10-CM classifications, showing that a better text embedding space trained by rich domain knowledge can greatly help visual representation learning.
Figure 3: ROC curves on Disorders and ICD-10-CM, including head/medium/tail parts respectively. The shadow in the figure shows the 95% CI (Confidence interval) and FM, KE is short for Fusion Module and Knowledge Enhancement. More ROC curves for each class are shown in Supplementary Section F